The perimeter of a semicircle is doubled when the radius is increased by 6.5. Find the radius of the semicircle.
Answer by mananth(16949)

let the radius be r
Perimeter of semicircle P = pi*r
..
increased radius = r+6.5
Perimeter = pi*(r+6.5)
pi*(r+6.5)= 2*pi*r
divide by pi
pi*(r+6.5)/pi= 2*pi*r/pi
r+6.5 = 2*r
add -r to both sides
r-r+6.5=2r-r
r=6.5, the radius.
..
perimeter of semicircle = 6.5*pi
= 20.4